OpenStudy (anonymous):

Hi, checking my answers 7(w-2)= -14 w=0? Thank you

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

7(w-2)=-14 w-2=-14/7=-2 w=-2+2=0